# Object-X: Learning to Reconstruct Multi-Modal 3D Object Representations

### v1

@@ Abstract
Learning effective multi-modal 3D representations of objects is essential for numerous applications, such as augmented reality and robotics. Existing methods often rely on task-specific embeddings that are tailored either for semantic understanding or geometric reconstruction. As a result, these embeddings typically cannot be decoded into explicit geometry and simultaneously reused across tasks. In this paper, we propose Object-X, a versatile multi-modal object representation framework capable of encoding rich object embeddings (e.g., images, point cloud, text) and decoding them back into detailed geometric and visual reconstructions. Object-X operates by geometrically grounding the captured modalities in a 3D voxel grid and learning an unstructured embedding fusing the information from the voxels with the object attributes. The learned embedding enables 3D Gaussian Splatting-based object reconstruction, while also supporting a range of downstream tasks, including scene alignment, single-image 3D object reconstruction, and localization. Evaluations on two challenging real-world datasets demonstrate that Object-X produces high-fidelity novel-view synthesis comparable to standard 3D Gaussian Splatting, while significantly improving geometric accuracy. Moreover, Object-X achieves competitive performance with specialized methods in scene alignment and localization. Critically, our object-centric descriptors require 3-4 orders of magnitude less storage compared to traditional image- or point cloud-based approaches, establishing Object-X as a scalable and highly practical solution for multi-modal 3D scene representation. The code is available at https://github.com/gaiadilorenzo/object-x.
Figure 1: Object-X learns object-centric embeddings from an input object segmentation of a 3D scene reconstruction. The embeddings learned from multi-modal data (e.g., mesh, images, text descriptions) enable fast 3D Gaussian Splat reconstruction via a specifically trained decoder, and other downstream tasks operating directly in the latent space, such as localization and scene alignment. Object-X allows for representing the scene as a set of object descriptors without having to store storage-heavy representations like point clouds and image databases, while providing similiar functionalities.
@@ 3.1 Structured Latents from Multi-View Images
Let a set of object instances $\mathcal{O}$ be given, where each object $o=(\mathcal{P},\mathcal{I},\mathcal{M},\mathcal{A},\dots)\in\mathcal{O}$ is associated with a multi-modal input, such as multi-view images ($\mathcal{I}$), instance masks ($\mathcal{M}$), point clouds ($\mathcal{P}$) and other attributes ($\mathcal{A}$). For each object $o$, we extract an object-specific latent embedding $\mathbf{w}_{o}$. The scene representation is defined as $\mathcal{W}=\{\mathbf{w}_{o}\}_{o\in\mathcal{O}}$, enabling lightweight storage compared to dense point clouds or images. The goal is to learn mapping $\mathcal{O}\;\mapsto\;\mathcal{W}$, where $\mathcal{W}$ allows decoding into 3D objects. Here, we focus on modalities $\mathcal{P}$ and $\mathcal{I}$ to learn the reconstructable embedding. Other modalities will be discussed when training also for downstream applications in the next section.
@@
Structured Latent Representation (SLat). We convert $f$ into a structured latent representation $z=\{(z_{i},p_{i})\}_{i=1}^{L}$ via a 3D encoder $E$ as follows:
@@ 3.2 Compression to a Dense Embedding
Although $z$ is sparse relative to raw 3D data, it can still be large for high-resolution grids when we aim at capturing fine details. Moreover, downstream tasks favor a compact embedding for each object, rather than a collection of per-voxel vectors. Thus, we learn a mapping to compress $z$ into a fixed-size vector $\mathbf{w}\in\mathbb{R}^{d}$, where $d$ is small and independent of the size of the object at hand as $\mathbf{w}=f_{\text{U-3DGS}}(z)$, where $\mathbf{w}$ is an Unstructured 3D Gaussian Splatting (U-3DGS) embedding. We implement $f_{\text{U-3DGS}}$ using a 3D network (similar to a U-Net) that first organizes $z$ into a dense $N\times N\times N\times C$ tensor (with zero-filling for inactive voxels), then, downsamples and encodes it to $\mathbf{w}$.
@@
$\mathcal{L}_{\text{compress}}\;=\;\frac{1}{L}\sum_{i=1}^{L}\Bigl{[}M_{i}\,\lVert\hat{z}_{i}-z_{i}\rVert^{2}\;+\;\tfrac{1-M_{i}}{w}\lVert\hat{z}_{i}-z_{i}\rVert^{2}\Bigr{]},$
@@ 3.3 Decoding to 3D Gaussian Splats
$\Theta\;=\;\mathcal{D}_{\text{GS}}(\hat{z}),\quad\Theta\;=\;\Bigl{\{}\!(x_{i},s_{i},q_{i},\alpha_{i},c_{i})\!\Bigr{\}}_{i=1}^{N}.$ (1)
@@
Figure 3: The proposed Object-X learns per-object embeddings that are beneficial for a number of downstream tasks, besides object-wise 3DGS reconstruction, such as cross-modal visual localization [13] (via image-to-object matching), 3D scene aligment [21] (via object-to-object matching), and full-scene reconstruction by integrating per-object Gaussians primitives.
@@ 3.4 Learning Auxiliary Tasks
Auxiliary Encoder Training with Frozen Core: Initially, the pre-trained U-3DGS encoder and decoder (responsible for appearance and geometry) are kept frozen. The newly introduced encoders for the auxiliary modalities are trained. In this stage, the learning is guided by a combined loss:
@@
where $\mathbf{w}_{\text{concat}}$ is the full concatenated embedding (U-3DGS + auxiliary features). $\mathcal{L}_{\text{task}}$ is the task-specific loss (e.g., a contrastive loss for localization). $\mathcal{L}_{\text{recon}}$ is the reconstruction loss also used for pre-training the U-3DGS embeddings. Crucially, $\mathcal{L}_{\text{recon}}$ is applied using the U-3DGS decoder, which operates only on the corresponding U-3DGS portion ($\mathbf{w}_{\text{U-3DGS}}$) of $\mathbf{w}_{\text{concat}}$. This ensures that the new auxiliary features are learned in a way that remains compatible with, and does not corrupt, the reconstructive capabilities of the core U-3DGS representation. $\lambda_{\text{recon}}$ balances these two objectives.
Joint Fine-tuning: After the auxiliary encoders have been trained, all network components are unfrozen. This includes the U-3DGS encoder and decoder, as well as all auxiliary modality encoders. The entire ensemble is then fine-tuned end-to-end using the same combined loss $\mathcal{L}_{\text{aux}}$ (Eq 2).
@@ 4 Experiments
ScanNet. To evaluate generalization, we test on ScanNet [4] without training our model on it. Since ScanNet does not provide scene graph annotations, we apply SceneGraphFusion [28] on RGB-D sequences to generate 3D instance segmentations and object relationships (used for auxiliary tasks). Compared to 3RScan, ScanNet captures RGB-D sequences at a higher frame rate with minimal motion between consecutive frames. To ensure diverse viewpoints, we sample one image every 25 frames. We use 77 test scenes from the split defined in [13]. Scenes, where SceneGraphFusion fails to generate annotations, are excluded. As in 3RScan, objects without associated images are discarded. The test split, along with its annotations, will be publicly released.
@@
For the baselines, we select $k$ frames that maximize viewpoint diversity. To achieve this, we apply $k$-means clustering to the positions of the cameras observing the object. In 3RScan, we use the maximum number of frames supported by DepthSplat (i.e., 12), while in ScanNet, where fewer non-overlapping frames are available, we limit the selection to four frames per object.
@@
Results on 3RScan. The top part of Table 1 reports the results on 3RScan. Object-X achieves comparable novel view synthesis quality to the baselines, with an SSIM score closest to 3DGS, the second-best PSNR score, and the best LPIPS score. Object-X substantially outperforms all baselines in geometric accuracy. Our method improves geometric accuracy by a large margin of 46 percentage points compared to all baselines while also exhibiting good completeness. This demonstrates that the proposed U-3DGS embeddings effectively capture object geometry, accurately recovered by the Object-X decoder. We omit geometric results for DepthSplat [30], which failed to produce reasonable geometry. Our runtime is three orders of magnitude faster than methods relying on optimization. Moreover, our approach requires an order-of-magnitude less storage, as we only store a single embedding per object instead of numerous images or 3D Gaussians.
Results on ScanNet. The bottom part of Table 1 reports the results on ScanNet. Note that we did not train our model on this dataset and used the model trained on 3RScan. Even without training, we achieve the highest novel view synthesis scores compared to the baselines, being the closest to the reference 3DGS reconstruction. Our geometric accuracy significantly outperforms all baselines. Given the higher resolution images in this dataset compared to 3RScan, the optimization-based approaches are quite slow, even when using only 4 views. We are four orders of magnitude faster than 3DGS (4V) and 2DGS (4V) requiring $3$ ms to reconstruct an object on average.
2. Scene Reconstruction. Although the we do not explicitly perform scene reconstruction, we evaluate scene composition by integrating object-level reconstructions. Object-X composes the full scene by decoding embeddings for each object and rendering their splats jointly. While this method is highly efficient, artifacts can emerge due to missing object segmentations. We also show results for Object-X + Opt, where the initial scene is constructed via Object-X, and optimized using 3DGS on the same image set used in 3DGS (12V). This setup isolates the benefit of Object-X as a strong initialization while allowing to overcome the problems caused by missing objects.
@@
3. Single-Image Reconstruction. We further evaluate the flexibility of Object-X by reconstructing objects from a single RGB or RGB-D frame. Let us note that the Object-X was not explicitly trained for this task nor to infer unseen parts of an object. In the RGB-only scenario, we estimate monodepth using [6]. Using the object mask and depth map, we lift the object pixels in 3D, obtaining a point cloud. We then voxelize it following the process described in Sec. 3.1. Object-X is then applied to obtain the object embedding from this input which is then fed directly into our decoder. We compare our approach to 3DGS [27] which optimizes 3D Gaussian splats based on a single masked image. Results in Table 1a show that Object-X outperforms 3DGS in all metrics on RGB-D and RGB inputs while also achieving significantly better F1 scores, indicating more precise and reliable reconstructions. We note that while there is still room for improvement in this task, achieving improved results demonstrate the versatility of the learned object embeddings.
@@
Method Modalities Mean RR $\uparrow$ Hits@1 $\uparrow$ Hits@2 $\uparrow$ Hits@3 $\uparrow$ Hits@4 $\uparrow$ Hits@5 $\uparrow$
$\mathcal{P}$ $\mathcal{O}$ 3DSG
@@ 5 Conclusion
We introduced Object-X, a novel framework for learning compact, versatile, multi-modal object-centric embeddings. These unique embeddings are decodable into high-fidelity 3D Gaussian Splats for object reconstruction while also serving as potent descriptors for diverse downstream tasks, such as visual localization, single-image reconstruction, and 3D scene alignment, often without task-specific fine-tuning. Object-X achieves excellent geometric accuracy and novel-view synthesis, comparable or superior to specialized methods, while drastically reducing storage requirements by 3-4 orders of magnitude by obviating the need for raw sensor data. Our approach effectively bridges the gap between abstract learned object representations and detailed explicit 3D models, offering a scalable and practical solution for advanced 3D understanding. The code will be made public.

### v2

@@ Abstract
Learning effective multi-modal 3D representations of objects is essential for numerous applications, such as augmented reality and robotics. Existing methods often rely on task-specific embeddings that are tailored either for semantic understanding or geometric reconstruction. As a result, these embeddings typically cannot be decoded into explicit geometry and simultaneously reused across tasks. In this paper, we propose Object-X, a versatile multi-modal object representation framework capable of encoding rich object embeddings (e.g., images, point cloud, text) and decoding them back into detailed geometric and visual reconstructions. Object-X operates by geometrically grounding the captured modalities in a 3D voxel grid and learning an unstructured embedding fusing the information from the voxels with the object attributes. The learned embedding enables 3D Gaussian Splatting-based object reconstruction, while also supporting a range of downstream tasks, including scene alignment, single-image 3D object reconstruction, and localization. Evaluations on two challenging real-world datasets demonstrate that Object-X achieves high-fidelity novel-view synthesis comparable to standard 3D Gaussian Splatting, while significantly improving geometric accuracy. Moreover, Object-X achieves competitive performance with specialized methods in scene alignment and localization. Critically, our object-centric descriptors require 3-4 orders of magnitude less storage compared to traditional image- or point cloud-based approaches, establishing Object-X as a scalable and highly practical solution for multi-modal 3D scene representation. The code is available at https://github.com/gaiadilorenzo/object-x.
Figure 1: Object-X learns object-centric embeddings from an input object segmentation of a 3D scene reconstruction. The embeddings learned from multi-modal data (e.g., mesh, images, text descriptions) enable fast 3D Gaussian Splat reconstruction via a specifically trained decoder, and other downstream tasks operating directly in the latent space, such as localization and scene alignment. Object-X allows for representing the scene as a set of object descriptors without having to store storage-heavy representations like point clouds and image databases, while providing similar functionalities.
@@ 3.1 Structured Latents from Multi-View Images
Let a set of object instances $\mathcal{O}$ be given, where each object $o=(\mathcal{P},\mathcal{I},\mathcal{M},\mathcal{A},\dots)\in\mathcal{O}$ is associated with a multi-modal input, such as multi-view images ($\mathcal{I}$), instance masks ($\mathcal{M}$), point clouds ($\mathcal{P}$) and other attributes ($\mathcal{A}$). For each object $o$, we extract an object-specific latent embedding $\mathbf{w}_{o}$. The scene representation is defined as $\mathcal{W}=\{\mathbf{w}_{o}\}_{o\in\mathcal{O}}$, enabling lightweight storage compared to dense point clouds or images. The goal is to learn mapping $\mathcal{O}\;\mapsto\;\mathcal{W}$, where $\mathcal{W}$ allows decoding into 3D objects. Here, we focus on modalities $\mathcal{P}$ and $\mathcal{I}$ to learn the reconstructable embedding. Other modalities will be discussed when training also for downstream applications in the next section. We denote the resulting latent object embedding as $\mathbf{w}$ which is the central representation used throughout the paper.
@@
Structured Latent Representation (SLat). We convert $f$ into a structured latent representation $z=\{(z_{i},p_{i})\}_{i=1}^{L}$ via a 3D encoder $\mathcal{E}$ as follows:
@@ 3.2 Compression to a Dense Embedding
Although $z$ is sparse relative to raw 3D data, it can still be large at high resolutions. Capturing fine details requires dense grids, which remain memory intensive. Moreover, downstream tasks favor a compact embedding for each object, rather than a collection of per-voxel vectors. Thus, we learn a mapping to compress $z$ into a fixed-size vector $\mathbf{w}\in\mathbb{R}^{d}$, where $d$ is small and independent of the size of the object at hand as $\mathbf{w}=f_{\text{U-3DGS}}(z)$, where $\mathbf{w}$ is an Unstructured 3D Gaussian Splatting (U-3DGS) embedding. We implement $f_{\text{U-3DGS}}$ using a 3D network (similar to a U-Net) that first organizes $z$ into a dense $N\times N\times N\times C$ tensor (with zero-filling for inactive voxels), then, downsamples and encodes it to $\mathbf{w}$.
@@
$\mathcal{L}_{\text{compress}}\;=\;\frac{1}{N^{3}}\sum_{i=1}^{N^{3}}\Bigl[M_{i}\,\lVert\hat{z}_{i}-z_{i}\rVert^{2}\;+\;\tfrac{1-M_{i}}{w}\lVert\hat{z}_{i}-z_{i}\rVert^{2}\Bigr],$
@@ 3.3 Decoding to 3D Gaussian Splats
$\Theta\;=\;\mathcal{D}_{\text{GS}}(\hat{z}),\quad\Theta\;=\;\Bigl\{\!(x_{i},s_{i},q_{i},\alpha_{i},c_{i})\!\Bigr\}_{i=1}^{M}.$ (1)
@@
Figure 3: The proposed Object-X learns per-object embeddings that are beneficial for a number of downstream tasks, besides object-wise 3DGS reconstruction, such as cross-modal visual localization [14] (via image-to-object matching), 3D scene alignment [22] (via object-to-object matching), and full-scene reconstruction by integrating per-object Gaussians primitives.
@@ 3.4 Learning Auxiliary Tasks
Auxiliary Encoder Training with Frozen Core: Initially, the pre-trained U-3DGS encoder and decoder (responsible for appearance and geometry) are kept frozen. The newly introduced encoders for the auxiliary modalities are trained. In this stage, learning is guided only by the task-specific loss $\mathcal{L}_{\text{task}}(\mathbf{w}_{\text{concat}})$, where $\mathbf{w}_{\text{concat}}$ is the full concatenated embedding (U-3DGS + auxiliary features). For example, $\mathcal{L}_{\text{task}}$ may be a contrastive loss for localization. This ensures that the new auxiliary features are learned in a way that remains compatible with, and does not corrupt, the reconstructive capabilities of the core U-3DGS representation.
Joint Fine-tuning: After the auxiliary encoders have been trained, all network components are unfrozen. The entire ensemble is then fine-tuned end-to-end using a combined objective:
@@
where $\mathcal{L}_{\text{recon}}$ is the reconstruction loss also used for pre-training the U-3DGS embeddings, applied using the U-3DGS decoder on the corresponding $\mathbf{w}_{\text{U-3DGS}}$. $\lambda_{\text{recon}}$ balances these two objectives. This stage allows for mutual adaptation of all parts of the representation, further optimizing for both the specific auxiliary task and the foundational 3D reconstruction quality.
@@ 4 Experiments
ScanNet. To evaluate generalization, we test on ScanNet [4] without training our model on it. Since ScanNet does not provide scene graph annotations, we apply SceneGraphFusion [30] on RGB-D sequences to generate 3D instance segmentations and object relationships (used for auxiliary tasks). This allows us to assess robustness to errors in 3D instance segmentation in the practical setting. Compared to 3RScan, ScanNet captures RGB-D sequences at a higher frame rate with minimal motion between consecutive frames. To ensure diverse viewpoints, we sample one image every 25 frames. We use 77 test scenes from the split defined in [14]. Scenes, where SceneGraphFusion fails to generate annotations, are excluded. As in 3RScan, objects without associated images are discarded. The test split, along with its annotations, will be publicly released.
@@
For the baselines, we select $k$ frames that maximize viewpoint diversity by applying $k$-means clustering to the positions of the cameras observing the object. In 3RScan, we use the maximum number of frames supported by DepthSplat (i.e., 12), while in ScanNet, where fewer frames are available, we limit the selection to four views per object. To ensure fairness, 3DGS/2DGS are also evaluated under this sparse-view setting ($\leq 12$ views), matching the input regime of DepthSplat and reflecting the natural sparsity of 3RScan/ScanNet. DepthSplat itself is evaluated in its intended setup on unmasked scene-level images, with object masks applied only after reconstruction, ensuring a consistent and fair comparison.
@@
Results on 3RScan. The top part of Table 1 reports the results on 3RScan. Object-X achieves comparable novel view synthesis quality to the baselines, with an SSIM score closest to 3DGS, the second-best PSNR score, and the best LPIPS score. Object-X substantially outperforms all baselines in geometric accuracy. Our method improves geometric accuracy by a large margin of 46 percentage points compared to all baselines while also exhibiting good completeness. This demonstrates that the proposed U-3DGS embeddings effectively capture object geometry, accurately recovered by the Object-X decoder. We omit geometric results for DepthSplat [32], which failed to produce reasonable geometry. We attribute the higher geometric accuracy to the voxel-grounded latent representation, which spatially constrains and aligns the decoded Gaussians to surfaces, leading to improved consistency compared to unstructured Gaussian sets.
Our runtime is three orders of magnitude faster than methods relying on optimization. Moreover, our approach requires an order-of-magnitude less storage, as we only store a single embedding per object instead of numerous images or 3D Gaussians.
Results on ScanNet. The bottom part of Table 1 reports the results on ScanNet. Note that we did not train our model on this dataset and used the model trained on 3RScan. Even without training, we achieve the highest novel view synthesis scores compared to the baselines, being the closest to the reference 3DGS reconstruction. Our geometric accuracy significantly outperforms all baselines. Because ScanNet provides higher-resolution images than 3RScan, optimization-based approaches are considerably slower, even when using only 4 views. We are four orders of magnitude faster than 3DGS (4V) and 2DGS (4V) requiring $3$ ms to reconstruct an object on average.
2. Scene Reconstruction. Although we do not explicitly perform scene reconstruction, we evaluate scene composition by integrating object-level reconstructions. Object-X composes the full scene by decoding embeddings for each object and rendering their splats jointly. While this method is highly efficient, artifacts can emerge due to missing object segmentations. We also show results for Object-X + Opt, where the initial scene is constructed via Object-X, and optimized using 3DGS on the same image set used in 3DGS (12V). This setup isolates the benefit of Object-X as a strong initialization while allowing to overcome the problems caused by missing objects.
@@
While we note that full-scene reconstruction is not the primary focus of this work, we include these experiments to demonstrate the viability of composing object-level embeddings into larger-scale reconstructions, where Object-X provides competitive quality, the highest geometric accuracy, and serves as a fast initializer for subsequent refinement.
3. Single-Image Reconstruction. We further evaluate the flexibility of Object-X by reconstructing objects from a single RGB or RGB-D frame. Let us note that the Object-X was not explicitly trained for this task nor to infer unseen parts of an object. In the RGB-only scenario, we estimate monodepth using [6]. Using the object mask and depth map, we lift the object pixels in 3D, obtaining a point cloud. We then voxelize it following the process described in Sec. 3.1. Object-X is then applied to obtain the object embedding from this input which is then fed directly into our decoder. We compare our approach to 3DGS [28] which optimizes 3D Gaussian splats based on a single masked image. Results in Table 1a show that Object-X outperforms 3DGS in all metrics on RGB-D and RGB inputs while also achieving significantly better F1 scores, indicating more precise and reliable reconstructions. We note that while there is still room for improvement in this task, achieving improved results demonstrate the versatility of the learned object embeddings. A promising direction, for example, is the principled integration of monocular 3D priors [33, 36, 29] could further enhance geometric fidelity without sacrificing efficiency.
Additionally, we compare Object-X with the recent MIDI-3D [8], a diffusion-based generative model for single-image 3D reconstruction. Unlike Object-X, which encodes metrically grounded geometry from observed data, MIDI is a purely generative approach that learns to synthesize plausible 3D shapes directly from RGB inputs. The results are shown in Table 2. While MIDI achieves impressive visual quality on in-distribution images similar to its training data, its performance drops significantly on out-of-distribution scenes, such as those in 3RScan, where object shapes, scales, and textures differ from common internet imagery. In these cases, MIDI often produces geometrically inconsistent outputs that remain visually plausible but lack realism.
In contrast, Object-X focuses on reconstruction rather than generation, leveraging voxel-grounded latent representations to maintain geometric consistency even under large appearance or domain shifts. As a result, it generalizes better to real-world scenes without requiring distribution-specific fine-tuning, producing reconstructions that are both structurally coherent and metrically faithful.
Method Accuracy@0.05 $\uparrow$ Completion@0.05 $\uparrow$ F1@0.05 $\uparrow$
MIDI [8] 29.522 44.342 34.516
Object-X (RGB) 43.480 57.214 48.397
Object-X (RGB-D) 65.780 61.759 63.223
Table 2: Comparison with MIDI [8] on a subset of 3RScan. Metrics are computed at a 5 cm threshold. Methods marked with * use automatic alignment without manual registration.
@@
Method Modalities Mean $R$ $\uparrow$ Hits@1 $\uparrow$ Hits@2 $\uparrow$ Hits@3 $\uparrow$ Hits@4 $\uparrow$ Hits@5 $\uparrow$
$\mathcal{P}$ $\mathcal{O}$ 3DGS
@@ 5 Conclusion
We introduced Object-X, a novel framework for learning compact, versatile, multi-modal object-centric embeddings. These unique embeddings are decodable into high-fidelity 3D Gaussian Splats for object reconstruction while also serving as potent descriptors for diverse downstream tasks, such as visual localization, single-image reconstruction, and 3D scene alignment, often without task-specific fine-tuning. Object-X achieves excellent geometric accuracy and novel-view synthesis, comparable or superior to specialized methods, while drastically reducing storage requirements by 3-4 orders of magnitude by obviating the need for raw sensor data. Our approach effectively bridges the gap between abstract learned object representations and detailed explicit 3D models, offering a scalable and practical solution for advanced 3D understanding.